OpenStudy (anonymous):

help please...ive gotten so far..and im stuck...a company will need 55,000 in 5 years for a new addition. to meet this goal the company deposits money in an account that pays 5% annual interest compound quarterly. find the amount that should be invested to total 55,000 in five years.

OpenStudy (reemii):

if you place a quantity A, how much to you get after one year? and after two years?

OpenStudy (reemii):

trick: A+ y percent of A = A (1+y) example: A+5% of A = A (1+ 0.05)

OpenStudy (anonymous):

ok let me show you how far i have gotten... i used a=p(1+r)^n 55000=p(1+0.0125)^20

OpenStudy (anonymous):

no i am stuck after that..

OpenStudy (reemii):

"5% annual rate compund quarterly" : doen't that just mean that it's %5% after one year?

OpenStudy (reemii):

typo: +5%

OpenStudy (anonymous):

wait is it 1.2820?

OpenStudy (reemii):

since we want to obtain 55,000 after 5 years, we can forget about the "quarterly compound" feature. So you look for the number p such that p (1+0.05)^5 = 55,000, right?

OpenStudy (anonymous):

i think im confused on how your doing it. dont i change the 5% to a decimal and thats how i got 0.0125

OpenStudy (reemii):

\(5\% = \frac{5}{100}\)